How can you please help me solve this loooooonnnnngggg example question.....please, please please! thank you! :)
Mr. Smith, a mathamethics teacher was travelig by bus with his family, consisting of his mother, wife, son and daughter. The bus conductor doubting the son's right to have a half fare ticket asked, "How old is your daughter?". Mr. Smith, irritated at having questioned...confused the conductor with his answer:
"My daughter is TWICE as old as my son. My wife is three times as old as the sum of the ages of my son and my daughter. And I am as old as my wife and my daughter. My Mother whose age is total of all our ages is now 69." Find the age of the daughter.

THANKS A LOT!.................:)

Answer by rapaljer(4551)

Let x = age of the son
2x= age of daughter
wife = 3(son + daugter)
wife = 3(x + 2x) = 3(3x) = 9x
math teacher - wife + daughter
math teacher = 9x + 2x = 11x
Total age = 69
Son + Daughter + wife + teacher = 69
x + 2x + 9x + 11x = 69
23x = 69
x=3 Son
2x = 6 Daughter
9x = 27 Wife
11x = 33 Teacher

Check: 3+6+27+33 = 69

R^2 at SCC